OpenStudy (anonymous):

What is the sum of the solutions of 2 lx-1l -4=-2

OpenStudy (jhannybean):

eliminate |x-1| by itself first. \[2|x-1|-4=-2\]\[2|x-1|=2\]\[|x-1|=1\] now |x-1| will have two solutions. -(x-1) and (x-1) \[-(x-1)=1 \]\[(x-1)=1\] solving the first one \[(x-1)=-1\]\[x-1+1=-1+1 \to x=0\] The second one \[x-1=1 \to x-1+1=1+1 \to x=2\]

OpenStudy (jhannybean):

so now you have two solutions, \[x=0 , x=2\] what is 2+0 ?

OpenStudy (anonymous):

2 right

OpenStudy (jhannybean):

Yup :)

OpenStudy (anonymous):

thank you very much :D

OpenStudy (jhannybean):

no problem :)